OpenStudy (anonymous):

Jack was finding the special product of a perfect square binomial. Check Jack’s work and identify the mistake that he made, and find the correct product. (4x−1) 2 = (4x) 2−2(4x) (1)+ (1) 2 =4x2−8x+1

OpenStudy (jadedry):

As you mentioned a perfect square binomial, I assume you meant this: \[(4x -1) ^ 2\] \[= (4x -1)(4x - 1)\] You have to deal with this in bite-size pieces First: \[4x * (4x -1)\] = \[(4x * 4x) + (4x * (-1))\] = \[16x^2 - 4x\] Multiply the right side by the (-1) in the first equation \[-1 * 4x + (-1) * (-1)\] =\[-4x + 1\] \[16x^2 - 4x + (-4x +1)\] = \[16x^2 - 8x + 1\] = Final product He squared x in 4x^2 , but neglected to square the coefficient.